OpenStudy (anonymous):

Find dy/dx y=cos^7(sin3x)

OpenStudy (danjs):

hi

OpenStudy (danjs):

\[\frac{ dy }{ dx }\cos ^{7}(\sin(3x)) = \]

OpenStudy (danjs):

right?

OpenStudy (danjs):

\[\frac{ d }{ dx }\cos^{7}[\sin(3x)] = \frac{ d }{ dx } [\cos(\sin(3x))^7 ]\]

OpenStudy (danjs):

gonna be using the chain rule a couple times

OpenStudy (danjs):

\[7 * \cos(\sin(3x))^6 * \frac{ d }{ dx }[\cos(\sin(3x))]\]

OpenStudy (danjs):

\[7*\cos(\sin(3x))^6 * -\sin(\sin(3x)) * \frac{ d }{ dx }3x\]

OpenStudy (danjs):

\[-21*\cos(3x) * \sin[\sin(3x)] * [\cos(\sin(3x))]^6\]

OpenStudy (danjs):

Final answer